OpenStudy (anonymous):

Please Help! Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 1^2 + 4^2 + 7^2 + ... + (3n - 2)^2 = n(6n^2 -3n -1) /2 I got as far as proving that the formula holds true for n=1. I need help with the rest please!

OpenStudy (anonymous):

@sami-21

OpenStudy (anonymous):

replace \(n\) by \(k\) and assert that \[1^2 + 4^2 + 7^2 + ... + (3k - 2)^2 = \frac{k(6k^2 -3k -1)}{2})\] is true, then prove the formula for \(k+1\)

OpenStudy (anonymous):

Right I'm sorry I forgot to say I did that too! I am having trouble proving formula for k+1

OpenStudy (anonymous):

the \(k+1\) statement is \[1^2 + 4^2 + 7^2 + ... + (3k - 2)^2 +(3(k+1)-2)^2\] \[= \frac{(k+1)(6(k+1)^2 -3(k+1) -1)}{2}\]

OpenStudy (anonymous):

and you can claim "by induction" that the left hand side is equal to \[\frac{k(6k^2 -3k -1)}{2}+(3(k+1)-2)^2\] and it is really ugly algebra from here on in

OpenStudy (anonymous):

so from that last do I try to FOIL?

OpenStudy (anonymous):

multiply out, combine like terms etc etc it is a pain i can't do it here, but according to wolfram it is \[1/2 (6 k^3+15 k^2+11 k+2)\]

OpenStudy (anonymous):

here is the left hand side of the equal sign http://www.wolframalpha.com/input/?i= \frac{k%286k^2+-3k+-1%29}{2}%2B%283%28k%2B1%29-2%29^2

OpenStudy (anonymous):

here is the right hand side http://www.wolframalpha.com/input/?i=%3D \frac{%28k%2B1%29%286%28k%2B1%29^2+-3%28k%2B1%29+-1%29}{2}

OpenStudy (anonymous):

That made my head hurt... so what do I do after? because it looks nothing like (k+1)

OpenStudy (anonymous):

you can see that they are the same

OpenStudy (anonymous):

give me one sec

OpenStudy (anonymous):

copy and paste, the link will not work as it is

OpenStudy (anonymous):

at least you will know what you are working towards

OpenStudy (anonymous):

they are the same just simplified right? but I thought they both had to equal (k+1)

OpenStudy (anonymous):

oh hell no

OpenStudy (anonymous):

what you get on the right is the formula replacing \(k\) by \(k+1\)

OpenStudy (anonymous):

i just copied and pasted, then replaced all the \(k\)'s i saw by \(k+1\)

OpenStudy (anonymous):

the left hand side is the sum from 1 to \(k+1\) instead of the sum from one two k

OpenStudy (anonymous):

good luck with the algebra, it looks punishing, but it is just algebra, and now you know what you should get when you do it

OpenStudy (anonymous):

Ok thanks!

OpenStudy (anonymous):

Could you help me with another? Or should I post it separate?